Show that y=x+tanx satisfies the differential equation cos2x d2y/dx2-2y+2x=0

1
Expert's answer
2021-01-07T15:51:24-0500

"y = x + \tan{x}\\

y' = 1 + \sec^2{x}\\

y'' = 2\sec^2{x}\tan{x}\\

\begin{aligned}
\cos^2{x}\cdot y'' - 2y + 2x &= 2\cos^2{x}\sec^2{x}\tan{x} \\&- 2x - 2\tan{x} + 2x \\&= 2\tan{x} - 2x - 2\tan{x} + 2x &= 0
\end{aligned}"